Show that if a≡b(mod m) and c≡d(mod m), where a, b, c, d ϵ Z with m≥2, then a-c≡b-d(mod m).
1
Expert's answer
2020-10-19T17:56:00-0400

Given an integer "m>1", called a modulus, two integers are said to be congruent modulo "m" , if "m" is a divisor of their difference (i.e., if there is an integer "k" such that "a-b=km"). Congruence modulo "m" is denoted: "a\equiv b(\mod m)".


If "a\equiv b(\mod m)" and "c\equiv d(\mod m)", then there exist integers "k" and "s" such that "a-b=km" and "c-d=sm". Then "(a-b)-(c-d)=km-sm". It follows that "(a-c)-(b-d)=(k-s)m" where "k-s" is an integer as a difference of two integers. Therefore,

"a-c\equiv b-d(\mod m)".